OpenStudy (anonymous):

State the value of each expression. 7^0

OpenStudy (zarkon):

1

jhonyy9 (jhonyy9):

is 1

OpenStudy (amistre64):

\begin{array}l\color{red}{\normalsize\text{7}}\color{orange}{\normalsize\text{^}}\color{#9c9a2e}{\normalsize\text{0}}\color{green}{\normalsize\text{ }}\color{blue}{\normalsize\text{=}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{1}}\color{red}{\normalsize\text{}}\end{array}

jhonyy9 (jhonyy9):

because every numbers one exponent zero is one

jhonyy9 (jhonyy9):

on exponent zero - sorry

OpenStudy (anonymous):

ohh wow, duh of course. thank you!

OpenStudy (amistre64):

it comes from the definition of exponents b^m --- = b^(m-n) b^n 7^4 ---- = 7^(4-4) = 7^0 = 1 7^4